A small bomb explodes in 2 pieces. A first piece of 1.20 kg moves at 18.0 m / s to the left. What is the speed in m / s of the second piece with a mass of 1.60 kg?


1
Expert's answer
2021-01-13T11:38:34-0500

Let's choose the rightward as a positive direction. Then, applying the law of conservation of momentum, we have:


"p_i=p_f,""0=-m_1v_1+m_2v_2,""v_2=\dfrac{m_1v_1}{m_2}=\dfrac{1.20\ kg\cdot 18.0\ \dfrac{m}{s}}{1.60\ kg}=13.5\ \dfrac{m}{s}."


The sign plus means that the second piece of the exploded bomb moves to the right.

Answer:

"v_2=13.5\ \dfrac{m}{s}," to the right.